OpenStudy (anonymous):

. A 25.0 L balloon is filled with 2.75 mol of helium. If the helium leaks out at the rate of 0.105 mol/hr, what is the volume of the balloon after 12.0 hr? Assume the pressure and temperature don’t change

OpenStudy (cuanchi):

you can use the Avogadro's Law V1/n1=V2/n2, isolate V2= V1n2/n1 to calculate n2 you have to multiply the leaks rate time the hours, 0.105 mol/h x 12.0h and subtract this value from the n1.